a man has 7 friends in how many ways can he invite one or more of them to a party
 

Grade:10

1 Answers

Arun
4 years ago
Dear Atharv
 
 
He can invite one or more friends by inviting 1 friend, or 2 friends or 3 friends, or all the 7 friends.
1 friend can be selected out of 7 in 7C1
2 friends can be selected out of 7 in 7C2
3 friends can be selected out of 7 in 7C3
4 friends can be selected out of 7 in 7C4
5 friends can be selected out of 7 in 7C5
6 friends can be selected out of 7 in 7C6
7 friends can be selected out of 7 in 7C7
Therefore the required number of ways (combinations) = 7C1 + 7C2 +7C3 + …..... + 7C7 = 27 – 1 = 128 – 1 = 127 ways
 
Regards
Arun (askIITians forum expert)